# Operating System And Artificial Intelligence: A Systematic Review

### v1

@@ 4.1 AI for OS Tuning
Wang et al. [87] introduce LearnedFTL,a page-level Flash Translation Layer design that harnesses the power of ML techniques to significantly enhance the random read performance of flash-based SSDs. This method effectively models the non-uniform yet largely linear relationships between logical page numbers and physical page numbers, overcoming the inherent irregularity in flash memory’s address space distribution.LearnedFTL embeds the training of learned indexes within the GC process, allowing for the continuous updating and refinement of the learned models as the address mapping evolves. LearnedFTL also incorporates a bitmap prediction filter acting as a safeguard against potential model inaccuracies, validating the predicted PPNs and ensuring that only correct mappings are used for address translation. This innovative approach uniquely minimizes the occurrence of double reads resulting from address translation during random read accesses.
Memory
Traditional heuristic-based methods struggle due to the intricate hierarchy of memory levels, dynamic workloads, data-dependent behavior, and hardware constraints. These complexities make it difficult to optimize memory management effectively.
Zhang et al.[17] address the challenges of address mapping in 3D-stacking memory, focusing on technologies like High-Bandwidth Memory and Hybrid Memory Cube. They propose Software-Defined Address Mapping, allowing user programs to directly control memory hardware. SDAM enables fine-grained data placement while leveraging chunk-based address mapping management. The system employs machine learning to identify access patterns, resulting in significant speedups compared to fixed address mapping systems.
Lagar-Cavilla et al.[88] introduce a software-defined far memory system that has been successfully deployed across Google’s data centers since 2016. The authors propose a proactive software-defined approach that compresses cold memory pages to create a far memory tier in software. This approach uses the Linux kernel’s zswap mechanism to compress pages and store them in DRAM, effectively creating a far memory tier with low latency. The authors designed an autotuning system that employs a Gaussian Process Bandit to estimate the size of cold memory and the promotion rate under different configurations. This model is used to emulate the control algorithm offline and estimate the impact of different parameter settings. The authors conclude that their software-defined far memory system is effective in saving memory costs without impacting application performance.
Rocha et al. [15] propose PredG, a Machine Learning framework designed to enhance the performance of graph processing. PredG aims to find the ideal thread and data mapping on Non-Uniform Memory Access systems. One of the key features of PredG is its ability to be agnostic to the input graph. It uses the available features of the graphs to train an Artificial Neural Network to perform predictions as new graphs arrive. This is done without any application execution after being trained, which makes it a powerful tool for optimizing graph execution on NUMA machines. This work is part of a broader trend in the research community towards automated graph machine learning, which seeks to discover the best hyper-parameter and neural architecture configuration for different graph tasks/data without manual design.
4.2 AI for OS security
With the widespread application of deep learning in recent years, using deep learning technologies for OS security has emerged, and the effectiveness of threat detection has been dramatically improved.
Qin et al. [35]presents MSNdroid, a novel malware detector designed specifically for Android applications, leverages a combination of native API calls, permissions, system API call features, and a Deep Belief Network. By applying deep learning techniques to native code features, MSNdroid effectively detects Android malware. This approach involves extracting features from a comprehensive dataset comprising malicious applications and benign applications. Notably, MSNDroid achieves an impressive accuracy while maintains an impressively low false-negative rate.
De Wit et al. [39] emphasize the value of incorporating machine learning in malware detection strategies for Android platforms. By leveraging accessible hardware data and sophisticated classification techniques, the study demonstrates the feasibility of identifying malware with a reasonable degree of precision, highlighting the potential of app-specific metrics in enhancing detection rates. This research contributes to the growing body of knowledge on AI-integrated security measures within operating systems, particularly pertinent to the Android ecosystem.
4.3 Findings
In response to Research Question 1, which explores the inclination of researchers to utilize AI in enhancing specific OS sub-domains, our comprehensive analysis of existing literature highlights two prominent areas: auto-tuning and security. These domains have attracted considerable attention due to their potential for transformative improvements through AI integration.
Auto-tuning emerges as a key area where AI is being employed to overcome the shortcomings of conventional heuristic-based methods. These traditional approaches often struggle to cope with the ever-evolving demands of modern computing environments. Researchers are harnessing machine learning to dynamically adjust OS parameters, such as scheduling, energy management, memory allocation, and storage optimization, aiming to boost system performance and resource efficiency. AI’s predictive abilities are particularly crucial in enhancing scheduling mechanisms to strike a balance between fairness, responsiveness, and throughput, effectively managing critical resources, ensuring they can handle the intricate and fluctuating workloads common in today’s computing landscapes.
The security domain has witnessed a surge in the adoption of deep learning technologies to combat malware threats, showcasing substantial advancements in detection efficacy. AI-powered malware detectors have proven superior in terms of accuracy and reduced false negatives compared to traditional machine learning models. Furthermore, AI frameworks are being developed to create adversarial malware, underscoring AI’s dual role in fortifying OS security while also exposing vulnerabilities in AI-based detection systems.
5 RQ2: How can OS be optimized to improve the efficiency of AI systems?
AI accelerators are different from traditional hardware, affecting all aspects of system design, from data-center scale to single-chip scale. They also add high requirement for system architecture, management, and programming[13].
Previous work has shown AI jobs critically demand high-speed I/O and low-latency and high-bandwidth data communication[18][21].Various attempt on hardware has been done to improve ML application performance,for exmaple,using newly NVMe SSD [89], relying on hardware FPGA for the I/O communication control instead of relying on OS-level interrupts that can significantly reduce both total I/O latency and its variance and algorithm level. Accessing hardware through the kernel introduces a performance bottleneck. To mitigate this bottleneck, one effective approach is to bypass the kernel altogether, enabling userspace programs to directly access hardware[20].
Bateni et al.[89] presents NeuOS, a comprehensive system solution aimed at providing latency predictability for executing multi-DNN workloads in autonomous systems while simultaneously managing energy optimization and dynamic accuracy adjustments. It coordinates system- and application-level solutions intelligently to ensure that multiple DNN instances operate efficiently and meet their respective deadlines to guarantee latency predictability. It manages energy consumption by dynamically adjusting parameters such as voltage and frequency scaling to minimize energy usage without compromising the latency predictability or accuracy requirements. Based on specific system constraints, NeuOS adjusts the accuracy level of DNN computations in real-time. This allows for trade-offs between computational precision and resource efficiency, ensuring that the system operates within its given constraints while maintaining an acceptable level of performance.
Wang et al.[21] showed what network for GPU AI remoting,a technique where the execution of GPU APIs is managed remotely through a network on a remote proxy instead of running GPU computations locally on the machine. The study takes a GPU-centric perspective to derive minimum latency and bandwidth requirements and aims to ensure that unmodified AI applications can run on remoting setups using commodity networking hardware without performance degradation. The paper introduces a novel theoretical framework that quantifies the minimum network requirements essential for efficient GPU API remoting. By formalizing the relationship between network latency, bandwidth, and remoting efficiency, this model provides foundational insights.
Serizawa et al. [19]proposed an solution focused on I/O bandwidth.The method aims to improve the reading performance of large training datasets by using high-performance I/O storage devices. The authors discusses the problem of copying datasets between local storage and shared storage and proposes a solution to conceal the time spent on copying by overlapping the copying and reading of training data.
5.1 Findings
Addressing Research Question 2, which delves into the ways OSes can elevate the efficiency of AI systems, our analysis underscores the pivotal function of OSes in handling the distinctive needs of AI tasks.
To optimize AI efficiency through the OS, several strategic approaches emerge as essential. Firstly, there is a need to tailor the OS to cater specifically to the requirements of AI accelerators. This includes facilitating high-speed I/O operations, minimizing latency in communications, and enabling autonomous operation of AI tasks. Secondly, the development of specialized runtime systems and schedulers becomes crucial, ensuring optimal allocation of resources and efficient execution of AI processes. Thirdly,optimizing I/O bandwidth further enhances the performance of AI applications. Collectively, these strategies form a comprehensive framework for improving the efficiency and effectiveness of AI systems through optimized OS integration.
6 RQ3: How and which AI techniques are typically used to improve OS?
To answer RQ3, we review all studies that investigate AI techniques applied in OS in this section. Among 38 AI4OS previous work,those did not clear clarify the AI tools excluded,we list the AI tools that are used more than twice. In general, RF, RNN (LSTM), KNN, RL, MLP,etc. are the most widely-used techniques.
6.1 Tools analysis

### v2

@@ 4.2 Storage systems
Another critical aspect of SSD performance is the efficiency of Flash Translation Layers (FTLs), which manage the mapping between logical and physical addresses. Wang et al. (?) propose LearnedFTL, a machine learning-enhanced page-level FTL design that tackles address translation inefficiencies. This method models the non-uniform yet largely linear relationships between logical and physical page numbers, overcoming the irregularity in flash memory’s address space distribution. LearnedFTL integrates the training of learned indexes within the garbage collection (GC) process, continuously updating and refining the models. A bitmap prediction filter safeguards against potential model inaccuracies, ensuring correct mappings for address translation, thus minimizing double reads during random read accesses. Compared to state-of-the-art FTLs, LearnedFTL reduces address-translation-induced double reads by up to 55.5%, achieving a P99 tail latency improvement of 2.9× to 12.2× and an average throughput improvement of 5.5× to 8.2× under random workloads.
Non-volatile memory (NVM) technologies face unique challenges, such as the write disturbance (WD) problem, which impacts reliability and endurance. Wu et al. (?) introduce LearnWD, a method that harnesses ML and out-of-place updates to mitigate the write disturbance (WD) problem in NVM. LearnWD employs clustering algorithms to categorize stale data based on their inherent error proneness. When a write request arrives, LearnWD analyzes the aggressor pattern of the new data and matches it to a stale block from the most suitable cluster, minimizing potential WD errors. By leveraging this speculative decision-making process, LearnWD reduces WD-prone writes while maintaining efficiency. Additionally, LearnWD requires only occasional retraining of its models to sustain its accuracy and performance. LearnWD is highly versatile and can be applied across various types of NVM, including single-level cell (SLC) and multi-level cell (MLC) configurations. Although the design is primarily optimized for phase-change memory (PCM), it is also extendable to other NVM technologies such as resistive RAM (ReRAM) and spin-transfer torque RAM (STT-RAM). Experimental results demonstrate that LearnWD achieves significant improvements: reducing WD errors by 20.1%, lowering write latency by 11.0%, and extending write endurance by 21.9% across 15 diverse real-world datasets.
Figure 3: General ML for Storage Enhancement Workflow
The aforementioned examples illustrate a typical pattern as shown in Figure.3, in which ML tools are often treated as a black box within storage systems to predict the future state of read/write requests. Once the decision-making process is complete, the operation that best aligns with the prediction is executed.
Insights and Challenges:
A significant trend in recent years is the integration of machine learning into storage systems, offering exciting opportunities to enhance performance and predictability in increasingly dynamic and complex environments. New type of storage hardware has stronger motivation for reserchers than those for schedulers. Unlike traditional heuristic-based methods, ML-driven solutions leverage data-driven models to capture non-linear relationships and adapt to diverse workloads in real-time. These approaches reduce manual intervention, improve resource utilization, and enable intelligent decision-making for tasks such as I/O scheduling, address translation, and error mitigation.
Despite these advancements, several key challenges remain. First, the overhead associated with inference in ML models can limit their practicality, particularly in high-concurrency or resource-constrained storage systems. Storage systems has higher time requirements. Second, current methods often target specific hardware components or workload patterns, making it difficult to generalize their applicability across different storage technologies and use cases. Third, ensuring the robustness and reliability of ML models under unpredictable conditions, such as workload spikes or hardware failures, remains an open problem.
4.3 Memory management
Memory management and storage management are fundamental components of an operating system, each responsible for optimizing resource utilization and providing abstraction for different hardware resources. Memory management focuses on volatile primary memory (RAM), abstracting physical memory into virtual memory for fast, temporary storage.
With the emergence of advanced hardware technologies such as 3D-stacked memory, significant improvements in memory bandwidth and latency have been realized. However, traditional heuristic-based memory address mapping algorithms often fail to fully leverage these advancements. To bridge this gap, Zhang et al. (?) propose Software-Defined Address Mapping (SDAM), a hardware-software co-design mechanism that dynamically adjusts address mapping strategies based on program access patterns, thereby maximizing hardware utilization in 3D memory. SDAM further integrates machine learning techniques to automatically identify and cluster access patterns of key program variables, enabling efficient and adaptive address mapping while minimizing overhead. This approach not only alleviates the burden on programmers but also allows the system to dynamically optimize memory performance for diverse workloads.
In parallel with the advancements in high-performance memory technologies, new memory paradigms such as far memory are reshaping how memory capacity is scaled in modern systems (?). Addressing similar challenges of dynamic workloads and cost efficiency, Lagar-Cavilla et al. (?) introduce a software-defined far memory system for warehouse-scale computers (WSCs) to address increasing memory demands and cost challenges. By leveraging zswap, a Linux kernel mechanism, the system proactively compresses and migrates cold memory pages to create a flexible, low-cost far memory tier in software. It employs machine learning-based autotuning to optimize performance under dynamic workloads while meeting strict service-level objectives (SLOs). In the online section, the kernel manages memory with two key modules: kstaled and kreclaimd. kstaled tracks page ages by scanning page table entries, while kreclaimd reclaims cold pages exceeding the Cold Age Threshold, moving them to zswap to free DRAM. The Node Agent bridges the kernel and offline components by collecting statistics, adjusting the Cold Age Threshold dynamically, and exporting data for further analysis. In the offline section, an ML-based Autotuner uses far memory data and a Gaussian Process Bandit algorithm to optimize parameter configurations for memory savings and performance. The tuned parameters are fed back to the Node Agent, enabling continuous system optimization. Successfully deployed at Google since 2016, the system achieves significant memory cost savings (4–5% TCO reduction) with minimal performance impact, demonstrating practicality and scalability for large-scale production environments.
Building on the theme of leveraging machine learning to optimize memory systems and workload performance, Rocha et al. (?) introduce PredG, a machine learning framework specifically designed to improve graph processing efficiency on Non-Uniform Memory Access (NUMA) systems. Unlike previous examples that focus on general-purpose optimization, PredG targets a specific application domain—graph processing. It identifies the optimal thread and data mappings for NUMA machines by utilizing graph features to train an artificial neural network for precise predictions. Notably, this framework eliminates the need for application execution after training, making it a powerful and efficient tool for optimizing graph workloads on NUMA architectures. This work aligns with the broader trend of automated graph machine learning, which aims to discover optimal hyper-parameters and neural architectures for diverse graph tasks and datasets without manual intervention.
New deployment paradigms are also emerging in memory management, reflecting the evolving landscape of modern computing systems. Virtualization, a critical component of contemporary computing environments, introduces unique challenges and opportunities for optimizing memory and resource management. Continuing the exploration of applying machine learning techniques to these areas, Miura et al. (?) tackle the issue of cache management in virtualized environments, which involve both guest and host operating systems. In such setups, the second cache (managed by the host OS) often experiences poor locality of reference, making traditional algorithms such as LRU less effective. To address this, the authors propose leveraging LSTM, a recurrent neural network model designed for time-series data, to enhance cache replacement decisions in the second cache. Through simulations and real-world workload evaluations, their approach demonstrates significant improvements in cache hit ratios and I/O throughput. This work highlights the potential of machine learning to address inefficiencies in traditional memory and storage systems, showcasing how advanced models can adapt to complex, dynamic environments. It aligns with the broader trend of using ML to optimize memory management, offering innovative solutions for improving performance in virtualized systems.
Insights and Challenges:
A significant trend in recent years is the emergence of new hardware and deployment paradigms, which have opened up exciting opportunities for researchers to explore innovative methods. Unlike traditional heuristic-based approaches, software-defined memory management frameworks, such as SDAM and software-defined far memory systems—leverage machine learning to dynamically adapt to workload behaviors. This not only reduces manual intervention but also improves resource utilization. Despite significant advancement in memory management, several key challenges remain. First, many existing approaches incur high inference overhead, which can limit their practicality, particularly in resource-constrained or latency-sensitive environments. Second, current methods often focus on specific aspects of optimization and rely heavily on domain expertise for design and tuning. This lack of generality highlights the need for more versatile and intelligent solutions that can adapt to a wide range of hardware and workload configurations. Future research should prioritize developing low-overhead, general-purpose methods that balance performance optimization with reduced reliance on specialized knowledge, enabling more efficient and scalable memory management frameworks.
4.4 Network
The rapid development of the Internet and mobile communication industries has significantly increased the complexity of modern network systems. This evolution is characterized by more intricate infrastructures, a greater diversity of connected devices and resources, and increasingly dynamic network structures (?). To meet the demands of this ever-changing environment, network systems must become more intelligent and agile, adapting to diverse device requirements and unpredictable conditions.
Artificial intelligence techniques, particularly deep neural networks (DNNs), have demonstrated significant potential in enhancing network traffic management. Kwon et al. (?) examine the deployment of DNNs for network traffic data classification, where DNNs are used to automatically classify real network traffic data collected from the ONOS (Open Network Operating System) platform. Experimental results with simple network topologies indicate that DNNs can be a promising approach to effective network packet classification. Additionally, it is confirmed that the deployment of DNNs for real network traffic data classification should consider both the data packets intended for delivery and those required for network maintenance, as the classification performance of DNNs significantly depends on the nature of the network traffic data.
Another critical aspect of modern networking is ensuring optimal data transmission rates under varying conditions. Li et al. (?) propose a solution named Machine Learning Rate Adaptation (MLRA), which uses ML techniques to identify correlations between transmission rates, throughput performance, and link quality, thereby achieving high-efficiency adaptability for the IEEE 802.11ac standard. MLRA employs a two-level rate search procedure and a congestion detection mechanism to ensure the scalability and thoroughness of the algorithm. Experimental results show that, on commercial IEEE 802.11ac NICs, MLRA outperforms existing RA solutions with performance improvements ranging from approximately 133.0% to 658.0%, with minimal overhead.
Beyond traffic and rate management, AI and Software-Defined Networking (SDN) are being combined to address broader challenges, such as network resilience during natural disasters. Veliyath et al. (?) introduce a novel approach to predicting, responding to, and rehabilitating from flood-based natural disasters using SDN, ML and image processing. It applies unsupervised KMeans clustering for optimal node distribution across satellite images and utilizes the supervised CatBoost algorithm with log transformation to forecast flood-affected areas. The study simulates an SDN controller using ONOS, known for its self-healing capabilities, next-generation device support, and high performance, and deploys a virtual network setup on ONOS using Mininet. Simulations show this approach can self-heal and enhance network performance.
Insights and Challenges:
The rapid evolution of mobile network systems, driven by increasingly complex infrastructure and dynamic network structures, demands intelligent and adaptive solutions for efficient management. The integration of AI and machine learning (ML) into network systems offers transformative potential. The researches highlights the potential of SDN with ML. SDN has ablitiy to flexibly adapt to different infrastructure and policies. Combining with ML tools, it demonstrate the versatility of AI in enhancing network resilience, efficiency and self-healing capabilities. These insights illustrate the growing role of AI-driven approaches in optimizing network performance, adaptability, and disaster recovery in increasingly dynamic and complex environments. Despite the promise of AI and ML in mobile network systems, several challenges remain. First, the performance of AI models, such as DNNs for traffic classification, heavily depends on the quality and diversity of the training data. The classification accuracy varies with the type of network traffic, requiring careful consideration of both delivery and maintenance packets. Second, while solutions like MLRA demonstrate significant performance gains, their scalability in highly dynamic and heterogeneous environments with diverse devices and traffic patterns is still a concern. Third, the computational overhead of deploying ML models, especially in real-time applications like traffic classification and disaster prediction, can strain resource-constrained network devices. Fourth, current mention methods still lack generalization ablitity across different platform and different targets. Addressing these challenges is crucial to fully realize the potential of AI in modern mobile network systems.
4.5 Ops
OS Ops are crucial for ensuring that systems function reliably and securely, thereby supporting the smooth operation of businesses and services that depend on them. Given the increasing complexity and scale of modern IT infrastructures, the potential for leveraging AI to improve OS Ops and maintenance is significant. Unlike the previously discussed sections, where ML modules are typically embedded within the OS to directly enhance scheduling, memory management, or storage systems, this section focuses on ML’s role as an external monitor. Here, ML operates outside the OS, analyzing system behaviors, logs, and workload patterns to make predictions and provide actionable insights.
A common starting point for such monitoring efforts is the analysis of system logs, which are the most accessible resources for engineers seeking clues about an OS’s status. Building on this idea, Dusane et al. (?) introduce LogEA, a specialized tool for forensic investigations of Linux-based systems. LogEA uses natural language processing (NLP) and machine learning to analyze system logs, identifying security breaches, unauthorized access, and other anomalies. By detecting log messages with negative sentiment, LogEA flags and alerts users to potential events of interest, offering a targeted approach to anomaly detection. This focus on log analysis highlights the importance of leveraging existing OS outputs to uncover irregularities.
Also based on logs, while LogEA focuses on forensic analysis and post-event investigation, Akram et al. (?) take a more proactive approach with Desh, a framework designed to diagnose and predict failures with short lead times. Desh applies a novel three-phase deep learning pipeline to logs, not only identifying chains of events that lead to failures but also predicting when and where failures are likely to occur. In the first phase, sequences of phrases leading to node failures are extracted, and LSTM Phase 1 is trained to identify and recognize such chains based on the training data. In the second phase, the chains formulated in Phase 1 are fed into LSTM Phase 2, which learns to account for the cumulative time differences ($\delta$ times) of the phrases within a failure chain, relative to the terminal phrase in that chain. During the inference phase, the chains learned in Phase 2 are used to estimate the lead times of future failures, along with predictions of the likely failure locations, based on test data that is completely independent of the training data. Desh utilizes an RNN with input/output layers and multiple hidden layers that form a stacked LSTM architecture, enabling effective training and prediction on system logs. Unlike LogEA, which alerts users to negative events retrospectively, Desh enables preemptive actions, such as migrating computation to healthy nodes, to minimize downtime. This shift from reactive forensics to proactive failure prediction showcases the versatility of ML in OS monitoring, as it can both analyze past events and anticipate future issues.
Android is the most widespread smartphone operating system, and its popularity has attracted attackers who develop various malicious applications. On the defensive front, considerable research has been dedicated to identifying the types and states of Android applications based on their system-level behavior. Maasmi et al. (?) use a ML technique to detect whether an application is running in the foreground or not. This technique aims to enhance the accuracy of behavioral malware detection by providing informative priors or contextual metadata on the app’s state to malware identification models. The results indicate that a structured ML pipeline, which first identifies the app and then detects its mode, can achieve substantially higher accuracy compared to direct mode identification.
In contrast to the log-based approaches above, Zhang et al. (?) propose a hardware-based solution for real-time workload forensics. Their method analyzes process execution flows by dividing them into consecutive frames and extracting Translation Lookaside Buffer (TLB) utilization features. These features are then fed into machine learning models to identify workloads at the granularity of individual processes. Unlike software-based workload forensics, which depend on OS services and can be compromised by software attacks, this hardware-based approach operates independently of the OS, offering a secure and efficient solution for monitoring workload behavior in real-time. This distinction highlights a critical difference: while tools like LogEA, Desh, and Maasmi’s pipeline leverage OS-level data for analysis, Zhang et al.’s method bypasses the OS entirely, focusing on hardware-level monitoring to enhance security and reliability.
Insights and Challenges:
Machine learning module as an external monitor for OS operations offers new opportunities to enhance reliability, security, and efficiency. Unlike embedded ML approaches, external monitoring uses logs, behaviors, and workload patterns to provide insights and predictions, enabling both forensic analysis and proactive failure prevention. However, challenges persist. Log-based methods depend on high-quality logs, and ensuring low overhead while maintaining real-time responsiveness can be difficult. Generalizability of ML models across different workloads and hardware configurations remains limited, requiring extensive retraining. Additionally, security and interpretability of ML solutions must be improved to prevent adversarial exploitation and ensure trust in critical systems. Monitoring is just one part of OS Ops. To truly reduce the burden on engineers, future AI-driven tools need to be more intelligent, general-purpose, and capable of self-healing, enabling systems to automatically detect, diagnose, and resolve issues without human intervention.
4.6 Security
OS security is one of the most prominent areas where ML tools are being employed to enhance functionality and robustness. ML-driven solutions are increasingly critical in addressing the evolving landscape of cyber threats, from malware detection to ransomware prevention and advanced persistent threats (APTs). These advancements leverage sophisticated algorithms to improve detection rates, reduce false positives, and provide proactive defense mechanisms across diverse platforms and attack types.
APTs are among the most challenging cybersecurity threats, characterized by their sophistication and ability to remain undetected for extended periods. APTs often evade traditional anomaly detection methods by exploiting system vulnerabilities and employing deceptive techniques. Detecting APTs with ML is particularly difficult due to the rarity of relevant datasets and the significant imbalance in data, which complicates accurate detection. To address this, Benabderrahmane et al. (?) introduce AE-APT, a novel method utilizing AutoEncoder (AE) architectures ranging from basic to Transformer-based designs. By incorporating online learning-based anomaly detection, AE-APT enables real-time analysis and immediate responses to ongoing APTs, offering a promising solution to one of the most complex areas of cybersecurity.
Malware detection is a key focus in AI-OS integration, particularly for Android, the most widely used mobile operating system. Qin et al. (?) introduce MSNdroid, an advanced malware detector that leverages deep learning techniques to analyze native API calls, permissions, system API features, and other application-level data. By utilizing a Deep Belief Network and a comprehensive dataset of malicious and benign applications, MSNdroid achieves high detection accuracy with an impressively low false-negative rate, making it a reliable tool for combating Android malware. Similarly, Alzaylaee et al. (?) propose DL-Droid, a deep learning-based system that employs dynamic analysis to detect malicious Android applications. By generating inputs based on application states, DL-Droid achieves a detection rate of up to 97.8% using dynamic features alone and 99.6% when combining dynamic and static features. This approach demonstrates the power of combining feature types to improve detection performance while enhancing code coverage. De Wit et al. (?) further explore malware detection on Android by incorporating hardware-level data and classification techniques, showing the potential of app-specific metrics to boost detection rates and contribute to the growing field of AI-enhanced OS security.
Ransomware remains one of the most critical cybersecurity threats, targeting systems worldwide and causing substantial financial and data losses. Wihar et al. (?) propose a novel approach to ransomware detection using an LSTM-based model capable of analyzing system call sequences in real-time. This model identifies anomalous behavior indicative of ransomware activities with high accuracy and low false-positive rates. Integrated into the Linux kernel, this method facilitates seamless communication and classification while maintaining efficient resource utilization. Experimental evaluations show that the LSTM model provides a proactive defense mechanism against ransomware threats, significantly enhancing the security framework of Linux systems through its adaptive learning capabilities.
Extending beyond detection, Jurečková et al. (?) propose a machine learning-based approach for the online clustering of malicious samples into malware families. Using a weighted k-nearest neighbor classifier for known families and an online k-means algorithm for emerging malware, their method achieves cluster purities ranging from 90.20% to 93.34%. This approach focuses on the static analysis of portable executable files for the Windows operating system, offering valuable insights for categorizing malware and understanding evolving threats.
Insights and Challenges:
Machine learning is transforming OS security by offering advanced capabilities in threat detection, behavior analysis, and adaptive defense. Unlike traditional rule-based systems, ML excels at identifying complex patterns in large, dynamic datasets, enabling the detection of previously unknown or evolving threats. By analyzing system behaviors, network traffic, and user activity, ML can not only identify anomalies but also predict potential attacks, allowing for proactive mitigation. Furthermore, ML’s ability to integrate static and dynamic features provides deeper insights into system vulnerabilities and attack surfaces, enhancing overall security frameworks. However, the deployment of ML in OS security reveals critical insights: robust models require diverse and high-quality training data, as well as mechanisms to adapt to the ever-changing nature of cyber threats. Real-time learning and decision-making are increasingly essential, as static models struggle to keep pace with sophisticated attackers. Moreover, the integration of ML into OS workflows must balance accuracy, speed, and resource efficiency to be practical in real-world environments. These advancements suggest that ML has the potential to shift OS security from reactive to proactive, but achieving this requires addressing challenges related to scalability, data scarcity, and adversarial robustness.
4.7 Analysis
4.7.1 Typical Pattern:
Figure 4 showed the general workflow for ML tools to assist OS.
Figure 4: General ML for OS Enhancement Workflow
The workflow for enhancing operating systems through machine learning is structured into two distinct phases: training and deployment. In the training phase, diverse datasets are collected from key system components, including the CPU, memory, storage, and I/O operations, serving as the foundation for model development. This raw data undergoes extensive preprocessing and transformation to ensure it is properly structured and optimized for training purposes. Advanced machine learning techniques are then employed to develop specialized models designed to address specific tasks, leveraging methods such as XGBoost, and various regression techniques.
In the deployment phase, the trained models are integrated into the operating system and applied to real-time operational data. By analyzing this incoming data, the system uses the predictive capabilities of the models to optimize performance dynamically. This phase involves fine-tuning configurations and generating actionable insights to improve key areas, including memory segment optimization, task prioritization, and dynamic resource reconfiguration. These enhancements illustrate the transformative potential of machine learning when systematically applied to operating systems, enabling more efficient, adaptive, and intelligent resource management.
4.7.2 How to choose proper algorithms to improve OS?
Among the previous studies surveyed, we figure out some popular algorithms, include Random Forests (RF/DT), Recurrent Neural Networks (RNN/LSTM), K-Nearest Neighbors (KNN), Reinforcement Learning (RL) and Multi-Layer Perceptrons (MLP).